OpenStudy (anonymous):

will medal and fan!!!! Raven’s parents bought a 20’ ski boat for $22,000. After 2 years, the boat was worth 17,500. Using x as the number of years and y as the boat value, graph the two points. Part 1: [2 points] Describe the slope of the line which connects the two points. Part 2: [3 points] How can your graph be used to predict the value of the boat in 5 years? Part 3: [2 points] What is the projected boat value in 5 years?

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

can u explaian to me this cuz i know this a lot

OpenStudy (anonymous):

*explain hoe to this

OpenStudy (the_fizicx99):

"Hoe" xD Yeah, I get what you mean, sure.

OpenStudy (anonymous):

lol oops *how

OpenStudy (anonymous):

xDDDD omg lol

OpenStudy (the_fizicx99):

(0, 22,000) and (2, 17,500) M = (y2 - y1) / (x2 - x1) \(\ M = \dfrac{ 17,500 - 22,000}{2 - 0} \implies \dfrac{-4,500}{2} \implies -2,250 \) y - y1 = m(x - x1) y - 22,000 = -2,250(x - 0) y - 22,000 = -2,250x +22,000 + 22,000 y = -2,250x + 22,000 Once you plot that, it'll be your function. y = f(x) so; f(5) = -2,250(5) + 22,000 --> -11,250 + 22,000 = 10,750 After 5 years, the boat will be worth $10,750

OpenStudy (anonymous):

0,o whoah THANK YOU SOOOOOOO MUCH lol *Hoe* do u know all this???

OpenStudy (anonymous):

lol

OpenStudy (the_fizicx99):

Oh, I studied this a long time ago ~ You can graph it, just enter "y = -2,250x + 22,000" into Desmos

OpenStudy (anonymous):

lol cool well thank u sooooo much really appreciated

OpenStudy (the_fizicx99):

For part B, it's asking for when you look at your graph, look for the x interval 5, and see what y interval corresponds to it. ^_^ Or you can use a function to make it easier

OpenStudy (the_fizicx99):

Yw